OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $70,000. It increased by the same factor each year and in 2006, the average annual salary was $82,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (anonymous):

f(x) = 70(1.17)x f(x) = 82(1.17)x f(x) = 70(2.2)x f(x) = 82(2.2)x

OpenStudy (anonymous):

It should be the second one.

OpenStudy (anonymous):

can you plz explain

OpenStudy (anonymous):

well we know the average annual salary for one year is 82000, and it hasn't changed so thats where the 82 comes from, and to get the annual factor simply divide 82000 by 70000, and you should get 1.17... and yea thats it really.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem.

OpenStudy (shes.radical):

Im 5 months late but shouldnt it be the first one..?